OpenStudy (anonymous):

Write (2 - 7i) in polar coordinates on the complex plane. (help)

OpenStudy (nurali):

2 - 7i x=2 and y=-7 r = √x^2+y^2 r = √(2^2 + (-7)^2) = √53 Tan ø=x/y Tan ø = 2/-7 = -2/7 arctan -2/7 = -74.05460409907715º ø = -74.05º

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

a+bi=r[cos(theta)+isin(theta)] do i have to put it in this form? ^

OpenStudy (anonymous):

Oh the Answer is (root(53), -74.05)

OpenStudy (anonymous):

Thank you:)

OpenStudy (nurali):

My Pleasure.